1.How long will it take an object to hit ground if it is thrown upward at 47 ft/sec from a height of 63 feet?

2.How do you know ahead of time how many roots a quadratic equation will have?

Answer by ankor@dixie-net.com(22740)

1.How long will it take an object to hit ground if it is thrown upward at 47 ft/sec from a height of 63 feet?
:
The height of an object thrown upward equation:
h = 0 when the object hits the ground
-16t^2 + 47t + 63 = 0
Solve this using the quadratic formula
a=-16; b=47; c=63
You should get a positive solution: x = 3.9375 sec to hit the ground
:
2.How do you know ahead of time how many roots a quadratic equation will have?
in the form y = ax^2 + bx + c, the discriminant:
D = -b^2 - 4*a*c
D > 0; two distinct real root
D = 0; one real number solution, a double root
D < 0; two complex number solutions